### v2

# Preserving the Hermiticity of the One-Body Density Matrix for a Non-Interacting Fermi Gas

Abstract
The one-body density matrix (ODM) for a $d$-dimensional zero temperature non-interacting Fermi gas can be approximately obtained in the semiclassical regime through different $\hbar$-expansion techniques. One would expect that each method of approximating the ODM should yield equivalent density matrices which are both Hermitian and idempotent to any order in $\hbar$. However, the Kirzhnits and Wigner-Kirkwood methods do not yield these properties, while the method of Grammaticos and Voros does. Here we show explicitly, for arbitrary $d\geq 1$-dimensions through an appropriate change into symmetric coordinates, that each method is indeed identical, Hermitian, and idempotent. This change of variables resolves the inconsistencies between the various methods, showing that the non-Hermitian and non-idempotent behaviour of the Kirzhnits and Wigner-Kirkwood methods is an artifact of performing a non-symmetric truncation to the semiclassical $\hbar$-expansions. Our work also provides the first explicit derivation of the $d\geq 1$-dimensional Grammaticos and Voros ODM, originally proposed by Redjati et al. [1] based on their $d=1,2,3,4$ expressions.
Keywords: Density-functional theory, semiclassical, $\hbar$-expansion, Hermitian, ultracold Fermi gas
1 Introduction
Density functional theory (DFT) is a powerful tool used to solve quantum many-body problems. It has a wide range of applications in different branches of science, see [2, 3, 4, 5, 6, 7] and references therein for some examples. Particularly, Hohenberg-Kohn-Sham DFT allows one to map $N$ interacting fermions to $N$ non-interacting fermions under the influence of an effective one-body potential $v_{\mathrm{eff}}(\mathbf{r})$ [8]. In $d\geq 1$-spatial dimensions the total-energy of the system can then be expressed as a unique functional of the one-body density matrix (ODM), $\rho$, by
$E[\rho]=T_{0}[\rho]+\int d^{d}rv_{\mathrm{eff}}(\mathbf{r})\rho(\mathbf{r}).$ (1.1)
$T_{0}$ is the kinetic energy functional, and $v_{\mathrm{eff}}[\rho]$ can be split into 1) a combined classical plus quantum interaction term $E_{\mathrm{int}}[\rho]$, and 2) a term accounting for the presence of any external potential $v_{\mathrm{ext}}(\mathbf{r})$:
$v_{\mathrm{eff}}(\mathbf{r})=\frac{\delta E_{\mathrm{int}}[\rho]}{\delta\rho}+v_{\mathrm{eff}}(\mathbf{r}).$ (1.2)
In practice Hohenberg-Kohn-Sham theory hinges on finding approximate expressions for the above density functionals, where perhaps the simplest is provided by the local-density approximation which assumes a simple form of $\rho$ that is locally valid even for inhomogeneous systems [9, 10]. Although this approximation is an excellent theoretical starting point it often fails to capture the complete physics of a given non-interacting Fermi gas since it neglects the effects of inhomogeneities. For non-uniform systems the exact ODM (and therefore exact density functionals) are in general not known, but density gradient corrections to the local-density approximation can be computed through various techniques [8, 10, 11, 12, 13]. A subset of these techniques include different semiclassical $\hbar$-expansions, three popular methods being the Kirzhnitz method (KODM) [14, 15, 16, 17], the Wigner-Kirkwood method (WKODM) [18], and the method of Grammaticos and Voros (GVODM) [19]. In reference [20] the KODM was derived up to second order in $\hbar$ for arbitrary $d\geq 1$-dimensions. Furthermore, in reference [10] the $d\geq 1$-dimensional WKODM method is discussed although the ODM is not explicitly derived, and in references [1, 21, 22] the GVODM is derived to second order in $\hbar$ for $d=1,2,3,4$-dimensions. Although in each case the ODMs were all derived for a zero temperature non-interacting Fermi gas in the presence of a position-dependent external potential, they can still be used to account for the effects of finite temperature and internal interactions, see references [22, 23].
The KODM and WKODM methods are Hermitian $*\rho(\mathbf{r},\mathbf{r}\prime)=\rho(\mathbf{r}\prime,\mathbf{r})^{*}$ and idempotent $\hat{\rho}^{2}=\hat{\rho}$ if all terms in the $\hbar$ expansion are accounted for, but they are not guaranteed to be if their expansions are truncated at a given order in $\hbar$. In particular, within reference [20] the $d\geq 1$-dimensional KODM is truncated to second order in $\hbar$, and is seen to be manifestly non-Hermitian and non-idempotent. In contrast, the GVODM preserves Hermiticity and idempotency regardless of whichever order in $\hbar$ the semiclassical expansion is truncated at. These issues are commented on in references [1, 21, 22]. The various ODMs should all be equivalent since they must yield the same physical properties of the Fermi gas, e.g., the total energy. Therefore it is unsatisfying that the GVODM is manifestly Hermitian and idempotent while the KODM and WKODM are not. Additionally, a $d\geq 1$-dimensional expression for the GVODM has not yet been explicitly derived. Although reference [1] proposed such an expression based on their $d=1,2,3,4$ results, a rigorous derivation was not performed.
The structure of this paper is as follows: In Section 2, we show that because the KODM is formulated in terms of inherently non-symmetric coordinates $(\mathbf{r},\mathbf{r}^{\prime})$ from the beginning of its derivation, and it does not display the same symmetry as the GVODM which is formulated in terms of center of mass $\mathbf{R}$ and relative $\mathbf{s}$ coordinates, shown in equation 1.3. These symmetric coordinates are typically better suited to help a system display any inherent symmetries it may possess, since the center of mass between two particles remains invariant under translations, rotations, and reflections of the entire system, while the distance between any two particles remains invariant under translations and rotations of the system with respect to $\mathbf{R}$.
$\mathbf{R}=(\mathbf{r}+\mathbf{r}^{\prime})/2,\hskip 56.9055pt\mathbf{s}=\mathbf{r}-\mathbf{r}^{\prime}.$ (1.3)
We explicitly show in $d\geq 1$-dimensions that, upon the proper symmetrization procedure (change of coordinates) from $(\mathbf{r},\mathbf{r}^{\prime})\rightarrow(\mathbf{R},\mathbf{s})$, the KODM is identical to the GVODM up to second order in $\hbar$. Our starting point relies on the ability to express each term of the KODM as a specific inverse Laplace transformation before the coordinate transformation is undertaken. We show in A that in terms of non-symmetric coordinates, the WKODM is identical to the KODM via these inverse Laplace transformations [10, 24], allowing us to make the necessary change of coordinates. In Section 3, we address the fact that the GVODM originates from the non-Hermitian and non-idempotent WKODM method in terms of non-symmetric coordinates, even though it results in a Hermitian and idempotent ODM in terms of symmetric coordinates. Similar to Section 2, we show in $d\geq 1$-dimensions that by applying a symmetric WKODM method in terms of $(\mathbf{R},\mathbf{s})$ coordinates from the very beginning, the resulting WKODM is identical to the GVODM to second order in $\hbar$ as expected.
The symmetrization procedure for both the KODM and WKODM demonstrates that regardless of whichever method is used to derive the ODM of a zero temperature non-interacting Fermi gas, the same Hermitian and idempotent expression is obtained as one would expect. Our results can be shown to hold when higher order terms in the ODMs are accounted for, but, for simplicity, we restricted our analysis to second order in $\hbar$. This therefore resolves the Hermiticity and idempotency inconsistency between the various methods, showing that problems only arise when non-symmetric truncations of semiclassical expansions are taken. Additionally this explicitly shows that the predicted $d\geq 1$-dimensional expression for the GVODM proposed in reference [1] is in fact correct, as we exactly obtain their result through our symmetrization procedure for both the KODM and WKODM methods. To our knowledge, these issues have been overlooked in the semiclassical expansion methods utilized in DFT.
2 Hermitian form of the KODM in $d\geq 1$-dimensions
The $d\geq 1$-dimensional KODM for a non-interacting Fermi gas is given in non-symmetric coordinates $(\mathbf{r},\mathbf{r}^{\prime})$ by equation 2.1 [20].
$\rho_{K}(\mathbf{r},\mathbf{r}^{\prime})=A_{0}+A_{1}+A_{2};$ (2.1)
$A_{0}$ $=\frac{gk_{F}^{d}}{(2\pi z)^{d/2}}J_{\frac{d}{2}}(z),$ (2.2)
$A_{1}$ $=-\frac{gk_{F}^{d-3}z^{2}}{4(2\pi z)^{d/2}}J_{\frac{d}{2}-1}(z)\bigg{(}\nabla_{\mathbf{r}}k_{F}^{2}\cdot\frac{\mathbf{s}}{s}\bigg{)},$ (2.3)
$A_{2}$ $=\frac{gk_{F}^{d-4}z^{2}}{24(2\pi z)^{d/2}}J_{\frac{d}{2}-2}(z)\big{(}\nabla_{\mathbf{r}}^{2}k_{F}^{2}\big{)}+\frac{gk_{F}^{d-6}z^{3}}{96(2\pi z)^{d/2}}J_{\frac{d}{2}-3}(z)\big{(}\nabla_{\mathbf{r}}k_{F}^{2}\big{)}^{2}$
$\quad+\frac{gk_{F}^{d-6}z^{4}}{32(2\pi z)^{d/2}}J_{\frac{d}{2}-2}(z)\bigg{(}\nabla_{\mathbf{r}}k_{F}^{2}\cdot\frac{\mathbf{s}}{s}\bigg{)}^{2}$
$\quad+\frac{gk_{F}^{d-4}z^{3}}{12(2\pi z)^{d/2}}J_{\frac{d}{2}-1}(z)\bigg{[}\nabla_{\mathbf{r}}\bigg{(}\nabla_{\mathbf{r}}k_{F}^{2}\cdot\frac{\mathbf{s}}{s}\bigg{)}\cdot\frac{\mathbf{s}}{s}\bigg{]}.$ (2.4)
The local Fermi wavenumber is defined by $k_{F}=k_{F}(\mathbf{r})=\frac{\sqrt{2m}}{\hbar}\big{(}\mu-V(\mathbf{r})\big{)}^{1/2}$ where, for convenience, we have suppressed the $\mathbf{r}$ coordinate dependence of both $k_{F}$ and the $A_{1,2,3}$ terms. $\mu$ is the chemical potential which is equal to the local Fermi energy for our case of a zero temperature Fermi gas, and $z=k_{F}s$ with $s=|s|=\sqrt{\mathbf{s}\cdot\mathbf{s}}$. $J_{\nu}(z)$ is a cylindrical Bessel function of order $\nu$ [25], and $g$ is a fermionic degeneracy factor where $g=1$ when the Fermi gas is spin polarized, and $g=2$ when it is not, since each fermion will then have two possible spin states.
Equation 2.1 was obtained by simplifying equation 17 from reference [20] using the Bessel function identities described in B, namely equations 2.1 and 2.2. Furthermore, we have collected the various terms of the ODM based on their order of $\hbar$ (or equivalently, their order of $\nabla_{\mathbf{r}}$). $A_{0}$ is the classical zeroth order in $\hbar$ term of the semiclassical expansion, also known as the Thomas-Fermi term. It is the only term which does not vanish in the local density regime, where effects due to inhomogeneities are sufficiently small such that $s$ is negligible compared to the scale over which the density changes, i.e., $s\approx 0$, yielding $\rho_{K}(\mathbf{r},\mathbf{r}^{\prime})\approx A_{0}$ [21, 22]. Going beyond the local density approximation and including quantum density gradient corrections yields $A_{1}$ and $A_{2}$, proportional to $\hbar$ and 2superscriptPlanck-constant-over-$\hbar^{2}$, respectively.
Equation 2.1 is manifestly non-Hermitian and non-idempotent as discussed in references [1, 21, 22]. However, by changing to center of mass $\mathbf{R}$ and relative coordinates $\mathbf{s}$ via equation 1.3, we show that equation 2.1 can be put into a manifestly Hermitian and idempotent form identical to the GVODM. The transformation from $(\mathbf{r},\mathbf{r}^{\prime})\rightarrow(\mathbf{R},\mathbf{s})$ variables is non-trivial as we will show. Consequently, to simplify the process we will individually transform each $A_{1,2,3}$ term in equation 2.1 and sum the results in the end via equation 2.1.
Although reference [20] derives equations 2.1, 2.2, 2.3 and 2.4 for $d\geq 1$-dimensions in terms of non-symmetric coordinates using the commutator method of Kirzhnitz [14], they can equivalently also be derived in these coordinates via the WKODM method [10, 18, 24]. We show this to be true in A so that equations 2.2, 2.3 and 2.4 may alternatively each be expressed by the following inverse Laplace transforms, respectively.
$A_{0}=\mathcal{L}^{-1}_{\mu}\bigg{\{}\frac{C_{0}(\mathbf{r},\mathbf{r}^{\prime};\beta)}{\beta}\bigg{\}},$ (2.5)
$A_{1}=\mathcal{L}^{-1}_{\mu}\bigg{\{}\frac{C_{1}(\mathbf{r},\mathbf{r}^{\prime};\beta)}{\beta}\bigg{\}},$ (2.6)
$A_{2}=\mathcal{L}^{-1}_{\mu}\bigg{\{}\frac{C_{2}(\mathbf{r},\mathbf{r}^{\prime};\beta)}{\beta}\bigg{\}},$ (2.7)
where $\beta$ is a complex parameter which is a conjugate variable to the chemical potential $\mu$, not to be confused with the inverse of the temperature $(k_{B}T)^{-1}$ [23]. $C(\mathbf{r},\mathbf{r}^{\prime};\beta)$ is known as the Bloch density matrix, which has been expanded to second order in $\hbar$ so that $C(\mathbf{r},\mathbf{r}^{\prime};\beta)\approx C_{0}(\mathbf{r},\mathbf{r}^{\prime};\beta)+C_{1}(\mathbf{r},\mathbf{r}^{\prime};\beta)+C_{2}(\mathbf{r},\mathbf{r}^{\prime};\beta)$. Similar to our descriptions of the $A_{1,2,3}$ terms, $C_{0}(\mathbf{r},\mathbf{r}^{\prime};\beta)$ is the zeroth order in $\hbar$ classical (Thomas-Fermi) term, $C_{1}(\mathbf{r},\mathbf{r}^{\prime};\beta)$ is the first order in $\hbar$ term, and $C_{2}(\mathbf{r},\mathbf{r}^{\prime};\beta)$ in the second order in $\hbar$ term. These are given by equations 2.8, 2.9 and 2.10 and the details their origin is addressed in more detail within A by equations 1.1, 1.2, 1.3, 1.4 and 1.5 [10].
$C_{0}(\mathbf{r},\mathbf{r}^{\prime};\beta)$ $=\frac{g}{(2\pi\hbar)^{d}}\int d^{d}p\;\textrm{e}^{-\beta H_{cl}(\mathbf{r},\mathbf{p})+\frac{i}{\hbar}\mathbf{p}\cdot\mathbf{s}},$ (2.8)
$C_{1}(\mathbf{r},\mathbf{r}^{\prime};\beta)$ $=-\frac{ig\hbar\beta^{2}}{2m(2\pi\hbar)^{d}}\int d^{d}p\;\textrm{e}^{-\beta H_{cl}(\mathbf{r},\mathbf{p})+\frac{i}{\hbar}\mathbf{p}\cdot\mathbf{s}}\Big{(}\nabla_{\mathbf{r}}V(\mathbf{r})\cdot\mathbf{p}\Big{)},$ (2.9)
$C_{2}(\mathbf{r},\mathbf{r}^{\prime};\beta)$ $=\frac{g\hbar^{2}}{(2\pi\hbar)^{d}}\int d^{d}p\;\textrm{e}^{-\beta H_{cl}(\mathbf{r},\mathbf{p})+\frac{i}{\hbar}\mathbf{p}\cdot\mathbf{s}}\bigg{(}-\frac{\beta^{3}}{4m}\nabla_{\mathbf{r}}^{2}V(\mathbf{r})+\frac{\beta^{3}}{6m}\big{(}\nabla_{\mathbf{r}}V(\mathbf{r})\big{)}^{2}$
$\hskip 99.58464pt-\frac{\beta^{4}}{8m^{2}}\big{(}\mathbf{p}\cdot\nabla_{\mathbf{r}}V(\mathbf{r})\big{)}^{2}+\frac{\beta^{3}}{6m^{2}}(\mathbf{p}\cdot\nabla_{\mathbf{r}})^{2}V(\mathbf{r})\bigg{)}.$ (2.10)
Here $H_{cl}(\mathbf{r},\mathbf{p})=\frac{\mathbf{p}^{2}}{2m}+V(\mathbf{r})$ is the classical counterpart of the quantum one-body Hamiltonian. Substituting equations 2.8, 2.9 and 2.10 into equations 2.5, 2.6 and 2.7 and applying the inverse Laplace transform relationship provided in B by equation 2.4 exactly yield equations 2.2, 2.3 and 2.4. Proof of this is explicitly shown in A.
This offers us the starting point for our symmetrization procedure where we will transform the variables of equations 2.8, 2.9 and 2.10 from $(\mathbf{r},\mathbf{r}^{\prime})\rightarrow(\mathbf{R},\mathbf{s})$ before taking the inverse Laplace transforms in equations 2.5, 2.6 and 2.7. Upon first glance one might attempt to directly transform the KODM [equations 2.1, 2.2, 2.3 and 2.4] into symmetric coordinates, however, by doing this one risks incorrectly neglecting the the appearance of additional terms of the symmetric semiclassical expansion proportional to both $\hbar$ and 2superscriptPlanck-constant-over-$\hbar^{2}$. These new terms form from the $\exp[-\beta V(\mathbf{r})]$ factor within the integrands in equations 2.8, 2.9 and 2.10 which we will show. For simplicity, we will perform the transformation of variables to each term separately, starting with $A_{0}$ and ending with $A_{2}$. The results will be summed afterwards via equation 2.1 to obtain the KODM in terms of the symmetric coordinates $(\mathbf{R},\mathbf{s})$.
2.1 $A_{0}$ change of variables
The $d\geq 1$-dimensional Gaussian integral in equation 2.8 can straightforwardly be evaluated to yield
$C_{0}(\mathbf{r},\mathbf{r}^{\prime};\beta)=g\bigg{(}\frac{m}{2\pi\beta\hbar^{2}}\bigg{)}^{d/2}\textrm{e}^{-\frac{m\mathbf{s}^{2}}{2\hbar^{2}\beta}-\beta V(\mathbf{r})}.$ (2.11)
To move into symmetric coordinates we use equation 1.3 to make the substitution $V(\mathbf{r})=V(\mathbf{R}+\mathbf{s}/2)$. Next we Taylor-expand this to second order in $\mathbf{s}$ (equivalently to second order in $\nabla_{\mathbf{R}}$) which yields equation 2.12. Since $\mathbf{s}\rightarrow 0$ is the limit of the local density approximation, and is equivalent to the limit where $\hbar\rightarrow 0$, this approximation is well justified.
$V(\mathbf{r})=V(\mathbf{R}+\mathbf{s}/2)\approx V(\mathbf{R})+\frac{1}{2}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}+\frac{1}{8}\Big{[}\nabla_{\mathbf{R}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}\cdot\mathbf{s}\Big{]}.$ (2.12)
Substituting equation 2.12 into equation 2.11, Taylor expanding the exponential with respect to small $\mathbf{s}$, and retaining only second order terms gives
$A_{0}\approx X_{a}+X_{b}+X_{c};$ (2.13)
$X_{a}$ $=g\bigg{(}\frac{m}{2\pi\hbar^{2}}\bigg{)}^{d/2}\mathcal{L}_{\mu}^{-1}\bigg{\{}\beta^{-d/2-1}\textrm{e}^{-\frac{m\mathbf{s}^{2}}{2\hbar^{2}\beta}-\beta V(\mathbf{R})}\bigg{\}},$ (2.14)
$X_{b}$ $=-g\bigg{(}\frac{m}{2\pi\hbar^{2}}\bigg{)}^{d/2}\bigg{(}\frac{1}{2}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}+\frac{1}{8}\Big{[}\nabla_{\mathbf{R}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}\cdot\mathbf{s}\Big{]}\bigg{)}$
×$\quad\times\mathcal{L}_{\mu}^{-1}\bigg{\{}\beta^{-d/2}\textrm{e}^{-\frac{m\mathbf{s}^{2}}{2\hbar^{2}\beta}-\beta V(\mathbf{R})}\bigg{\}},$ (2.15)
$X_{c}$ $=\frac{g}{8}\bigg{(}\frac{m}{2\pi\hbar^{2}}\bigg{)}^{d/2}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}^{2}\mathcal{L}_{\mu}^{-1}\bigg{\{}\beta^{-d/2+1}\textrm{e}^{-\frac{m\mathbf{s}^{2}}{2\hbar^{2}\beta}-\beta V(\mathbf{R})}\bigg{\}},$ (2.16)
where the $X_{a,b,c}$ terms have been separated based upon their power of $\beta$. $X_{a}$ originates from the zeroth order part of the exponential expansion, where $\exp[-\beta V(\mathbf{R}+\mathbf{s}/2)]$ is simply replaced by $\exp[-\beta V(\mathbf{R})]$. It is the same in form as equation 2.2 but with $\mathbf{r}$ replaced by $\mathbf{R}$. $X_{b}$, and $X_{c}$ come from the first and second order parts of the exponential expansion, only appearing when the coordinate change is performed before the inverse Laplace transform is taken. Each inverse Laplace transform in equations 2.14, 2.15 and 2.16 can be evaluated via equation 2.4 which is derived in B. After substituting the result and simplifying, the expression for $A_{0}$ in symmetric $(\mathbf{R},\mathbf{s})$ coordinates is obtained and given by
$A_{0}$ $=\frac{gk_{F}^{d}}{(2\pi z)^{d/2}}J_{\frac{d}{2}}(z)+\frac{gk_{F}^{d-3}z^{2}}{4(2\pi z)^{d/2}}J_{\frac{d}{2}-1}(z)\bigg{(}\nabla_{\mathbf{R}}k_{F}^{2}\cdot\frac{\mathbf{s}}{s}\bigg{)}$
$\quad+\frac{gk_{F}^{d-4}z^{3}}{16(2\pi z)^{d/2}}J_{\frac{d}{2}-1}(z)\bigg{[}\nabla_{\mathbf{R}}\bigg{(}\nabla_{\mathbf{R}}k_{F}^{2}\cdot\frac{\mathbf{s}}{s}\bigg{)}\cdot\frac{\mathbf{s}}{s}\bigg{]}$
$\quad+\frac{gk_{F}^{d-6}z^{4}}{32(2\pi z)^{d/2}}J_{\frac{d}{2}-2}(z)\bigg{(}\nabla_{\mathbf{R}}k_{F}^{2}\cdot\frac{\mathbf{s}}{s}\bigg{)}^{2}.$ (2.17)
We also used the definition $k_{F}=k_{F}(\mathbf{R})=\frac{\sqrt{2m}}{\hbar}\big{(}\mu-V(\mathbf{R})\big{)}^{1/2}$ so that $k_{F}$ is now a function of $\mathbf{R}$. This will also be the case for the symmetrized $A_{1}$ and $A_{2}$ terms derived later. The first term in equation 2.17 is precisely in the same form as equation 2.2, as expected, and the three new additional terms are first and second order gradient corrections similar in form to parts of equations 2.3 and 2.4.
2.2 $A_{1}$ change of variables
Before changing the coordinates of equation 2.9, it can be re-expressed by replacing the $\mathbf{p}$ factor with one proportional to $\nabla_{\mathbf{s}}$ and solving the resulting $d\geq 1$-dimensional Gaussian integral in a similar manner to the previous section. This gives
$C_{1}(\mathbf{r},\mathbf{r}^{\prime};\beta)$ $=-\frac{ig\hbar\beta^{2}}{2m(2\pi\hbar)^{d}}\int d^{d}p\;\textrm{e}^{-\beta H_{cl}(\mathbf{r},\mathbf{p})+\frac{i}{\hbar}\mathbf{p}\cdot\mathbf{s}}\Big{(}\nabla_{\mathbf{r}}V(\mathbf{r})\cdot\mathbf{p}\Big{)}$
$=-\frac{g\hbar^{2}\beta^{2}}{2m(2\pi\hbar)^{d}}\Big{(}\nabla_{\mathbf{r}}V(\mathbf{r})\cdot\nabla_{\mathbf{s}}\Big{)}\int d^{d}p\;\textrm{e}^{-\beta H_{cl}(\mathbf{r},\mathbf{p})+\frac{i}{\hbar}\mathbf{p}\cdot\mathbf{s}}$
$=-\frac{gm^{d/2-1}\hbar^{2-d}}{2(2\pi)^{d/2}}\Big{(}\nabla_{\mathbf{r}}V(\mathbf{r})\cdot\nabla_{\mathbf{s}}\Big{)}\beta^{-d/2+2}\textrm{e}^{-\frac{m\mathbf{s}^{2}}{2\hbar^{2}\beta}-\beta V(\mathbf{r})}.$ (2.18)
For symmetrization, we again approximate $V(\mathbf{r})$ to second order using equation 2.12. This time we also make use of the approximate gradient identity derived in B, equation 2.8, to transform $\nabla_{\mathbf{r}}V(\mathbf{r})$ into $(\mathbf{R},\mathbf{s})$ coordinates. Combining this with equations 2.18 and 2.6 yields
$A_{1}\approx Y_{a}+Y_{b}+Y_{c};$ (2.19)
$Y_{a}$ $=-\frac{gm^{d/2-1}\hbar^{2-d}}{2(2\pi)^{d/2}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\nabla_{\mathbf{s}}\big{)}\mathcal{L}_{\mu}^{-1}\bigg{\{}\beta^{-d/2+1}\textrm{e}^{-\frac{m\mathbf{s}^{2}}{2\hbar^{2}\beta}-\beta V(\mathbf{R})}\bigg{\}},$ (2.20)
$Y_{b}$ $=\frac{gm^{d/2-1}\hbar^{2-d}}{4(2\pi)^{d/2}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\nabla_{\mathbf{s}}\big{)}\mathcal{L}_{\mu}^{-1}\bigg{\{}\beta^{-d/2+2}\textrm{e}^{-\frac{m\mathbf{s}^{2}}{2\hbar^{2}\beta}-\beta V(\mathbf{R})}\bigg{\}},$ (2.21)
$Y_{c}$ $=-\frac{gm^{d/2}\hbar^{2-d}}{4(2\pi)^{d/2}}\Big{[}\nabla_{\mathbf{R}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}\cdot\nabla_{\mathbf{s}}\Big{]}\mathcal{L}_{\mu}^{-1}\bigg{\{}\beta^{-d/2+1}\textrm{e}^{-\frac{m\mathbf{s}^{2}}{2\hbar^{2}\beta}-\beta V(\mathbf{R})}\bigg{\}}.$ (2.22)
Here we have not separated each term based on it’s power of $\beta$ like we did for equations 2.14, 2.15 and 2.16, and instead separate them based on $\nabla_{\mathbf{R}}$. Using the inverse Laplace transformation identity equation 2.4 as before and the gradient and Bessel function relationships of equations 2.7 and 2.3 from B, we can evaluate and simplify equations 2.19, 2.21 and 2.22 to find $A_{1}$ in terms of symmetric coordinates
$A_{1}$ $=-\frac{gk_{F}^{d-3}z^{2}}{4(2\pi z)^{d/2}}J_{\frac{d}{2}-1}(z)\bigg{(}\nabla_{\mathbf{R}}k_{F}^{2}\cdot\frac{\mathbf{s}}{s}\bigg{)}$
$\quad-\frac{gk_{F}^{d-6}z^{4}}{16(2\pi z)^{d/2}}J_{\frac{d}{2}-2}(z)\bigg{(}\nabla_{\mathbf{R}}k_{F}^{2}\cdot\frac{\mathbf{s}}{s}\bigg{)}^{2}$
$\quad-\frac{gk_{F}^{d-4}z^{3}}{8(2\pi z)^{d/2}}J_{\frac{d}{2}-1}(z)\bigg{[}\nabla_{\mathbf{R}}\bigg{(}\nabla_{\mathbf{R}}k_{F}^{2}\cdot\frac{\mathbf{s}}{s}\bigg{)}\cdot\frac{\mathbf{s}}{s}\bigg{]}.$ (2.23)
The first term of equation 2.23 is in the same form as equation 2.3 since equation 2.20 comes from the zeroth order part of the expansion of $\exp[-\beta V(\mathbf{r})]$ in equation 2.9, while the additional terms $Y_{b}$ and $Y_{c}$ yield second order gradient corrections. This is similar to the the comments we made regarding equation 2.17.
2.3 $A_{2}$ change of variables
To symmetrizize the final term $A_{2}$, we take a different approach which avoids the need to perform the coordinate transformation before the inverse Laplace transform is evaluated. Since we are only keeping up to second order terms and every term within the integrand of equation 2.10 is already second order in $\nabla_{\mathbf{r}}$, when expanding $\exp[-\beta V(\mathbf{r})]$ using equation 2.12, only the zeroth order part $\exp[-\beta V(\mathbf{R})]$ is retained. Therefore, no additional second order or lower terms appear in the symmetric expression for $A_{2}$. This is in contrast to $A_{0}$ and $A_{1}$ where the additional terms $X_{b,c}$ and $Y_{b,c}$ appeared, respectively. Thus for $A_{2}$, we can simply directly apply our transformation of variables equation 1.3 to the original non-symmetric KODM expression of $A_{2}$, equation 2.4. To do this, we use three gradient identities given by equations 2.8, 2.9 and 2.10 as shown and derived in B. Substituting the identities into equation 2.4 yields the symmetric value of $A_{2}$, which after simplification is given by
$A_{2}$ $\approx\frac{gk_{F}^{d-4}z^{2}}{24(2\pi z)^{d/2}}J_{\frac{d}{2}-2}(z)\big{(}\nabla_{\mathbf{R}}^{2}k_{F}^{2}\big{)}+\frac{gk_{F}^{d-6}z^{3}}{96(2\pi z)^{d/2}}J_{\frac{d}{2}-3}(z)\big{(}\nabla_{\mathbf{R}}k_{F}^{2}\big{)}^{2}$
$\quad+\frac{gk_{F}^{d-6}z^{4}}{32(2\pi z)^{d/2}}J_{\frac{d}{2}-2}(z)\bigg{(}\nabla_{\mathbf{R}}k_{F}^{2}\cdot\frac{\mathbf{s}}{s}\bigg{)}^{2}$
$\quad+\frac{gk_{F}^{d-4}z^{3}}{12(2\pi z)^{d/2}}J_{\frac{d}{2}-1}(z)\bigg{[}\nabla_{\mathbf{R}}\bigg{(}\nabla_{\mathbf{R}}k_{F}^{2}\cdot\frac{\mathbf{s}}{s}\bigg{)}\cdot\frac{\mathbf{s}}{s}\bigg{]}.$ (2.24)
As expected, this is exactly in the same form as equation 2.4, but with $\mathbf{r}$ replaced by $\mathbf{R}$, similar to how the first terms in equations 2.17 and 2.23 are in the same form as equations 2.2 and 2.3, respectively.
2.4 Symmetrized KODM
We are finally in a position to obtain the symmetric KODM by combining the symmetrized $A_{1,2,3}$ terms. By substituting equations 2.17, 2.23 and 2.24 into equation 2.1, and after simplification, we obtain equation 2.25 which is the key result of this paper. As a reminder, $k_{F}=k_{F}(\mathbf{R})=\frac{\sqrt{2m}}{\hbar}\big{(}\mu-V(\mathbf{R})\big{)}^{1/2}$.
$\rho_{K}(\mathbf{R},\mathbf{s})$ $=\frac{gk_{F}^{d}}{(2\pi z)^{d/2}}J_{\frac{d}{2}}(z)+\frac{gk_{F}^{d-4}z^{2}}{24(2\pi z)^{d/2}}J_{\frac{d}{2}-2}(z)\big{(}\nabla_{\mathbf{R}}^{2}k_{F}\big{)}$
$\quad+\frac{gk_{F}^{d-6}z^{3}}{96(2\pi z)^{d/2}}J_{\frac{d}{2}-3}(z)\big{(}\nabla_{\mathbf{R}}k_{F}\big{)}^{2}$
$\quad+\frac{gk_{F}^{d-4}z^{3}}{48(2\pi z)^{d/2}}J_{\frac{d}{2}-1}\bigg{[}\nabla_{\mathbf{R}}\bigg{(}\nabla_{\mathbf{R}}k_{F}\cdot\frac{\mathbf{s}}{s}\bigg{)}\cdot\frac{\mathbf{s}}{s}\bigg{]}.$ (2.25)
Equation 2.25 is exactly identical to the $d\geq 1$-dimensional GVODM originally proposed by reference [1] based upon their $d=1,2,3,4$ derivations. Here we have provided an explicit derivation of their result starting from the known $d\geq 1$-dimensional KODM expression of reference [20], and have shown that upon the proper change to symmetric coordinates, the KODM and GVODM are identical, $\rho_{K}(\mathbf{R},\mathbf{s})=\rho_{GV}(\mathbf{R},\mathbf{s})$. The loss of both Hermiticity and idempotency of the KODM is only an artifact of a non-symmetric truncation of the semiclassical $\hbar$ expansion.
3 Hermitian form of the WKODM in $d\geq 1$-dimensions
3.1 Symmetric Bloch matrix
In references [1, 21, 22], the manifestly Hermitian and idempotent GVODM in symmetric coordinates is derived for $d=1,2,3,4$-dimensions despite the fact that it is through means of a non-Hermitian and non-idempotent WKODM method in terms of non-symmetric coordinates. In this section, we will provide a $d\geq 1$-dimensional derivation completely in terms of symmetric coordinates starting from the same point that reference [10] uses for their non-symmetric WKODM derivation, see their section 4.3 where the non-symmetric Bloch matrix is given by
$C(\mathbf{r},\mathbf{r}^{\prime};\beta)=g\langle\mathbf{r}|\textrm{e}^{-\beta\hat{H}}|\mathbf{r}^{\prime}\rangle.$ (3.1)
Note that one can easily derive the presence of the degeneracy factor $g$ by accounting for fermionic spin within the position states and simplifying. $\hat{H}$ is the quantum one-body Hamiltonian, and we also note that the WKODM approximation has yet to be applied. Before changing to the symmetric coordinates, consider the unitary translation operator $\hat{U}(\mathbf{a})=\exp[-(i/\hbar)\hat{\mathbf{P}}\cdot\mathbf{a}]$, where $\hat{\mathbf{P}}=-i\hbar\nabla_{\mathbf{R}}$ is the $\mathbf{R}$-space representation of the momentum operator $\hat{\mathbf{P}}|\mathbf{p}\rangle=\mathbf{p}|\mathbf{p}\rangle$, and $\mathbf{a}$ is an arbitrary position space vector. $\hat{U}(\mathbf{a})$ acts on position space states and scalar functions $f$ as follows [26]
$\hat{U}(\mathbf{a})|\mathbf{R}\rangle$ $=|\mathbf{R}-\mathbf{a}\rangle,$
$\hat{U}(\mathbf{a})^{\dagger}|\mathbf{R}\rangle$ $=|\mathbf{R}+\mathbf{a}\rangle,$
$\hat{U}(\mathbf{a})^{\dagger}f(\mathbf{b})\hat{U}(\mathbf{a})$ $=f(\mathbf{a}+\mathbf{b}),$
where $\mathbf{b}$ is an arbitrary position space vector. Applying equation 1.3 to equation 3.1 and using the above unitary identities gives the expression for the Bloch matrix in terms of symmetric coordinates by
$C(\mathbf{R},\mathbf{s};\beta)$ $=g\langle\mathbf{R}+\mathbf{s}/2|\textrm{e}^{-\beta\hat{H}}|\mathbf{R}-\mathbf{s}/2\rangle$
$=g\langle\mathbf{R}|\textrm{e}^{-\frac{i}{\hbar}\hat{\mathbf{P}}\cdot\mathbf{s}/2}\textrm{e}^{-\beta\hat{H}}\textrm{e}^{-\frac{i}{\hbar}\hat{\mathbf{P}}\cdot\mathbf{s}/2}|\mathbf{R}\rangle$
$=g\langle\mathbf{R}|\textrm{e}^{-\frac{i}{\hbar}\hat{\mathbf{P}}\cdot\mathbf{s}}\hat{U}(\mathbf{s}/2)^{\dagger}\textrm{e}^{-\beta\hat{H}}\hat{U}(\mathbf{s}/2)|\mathbf{R}\rangle$
$=g\langle\mathbf{R}|\textrm{e}^{-\frac{i}{\hbar}\hat{\mathbf{P}}\cdot\mathbf{s}}\textrm{e}^{-\beta\hat{U}(\mathbf{s}/2)^{\dagger}\hat{H}\hat{U}(\mathbf{s}/2)}|\mathbf{R}\rangle$
$=g\langle\mathbf{R}|\textrm{e}^{-\frac{i}{\hbar}\hat{\mathbf{P}}\cdot\mathbf{s}}\textrm{e}^{-\beta\frac{\hat{\mathbf{P}}^{2}}{2m}-\beta V(\mathbf{R}+\mathbf{s}/2)}|\mathbf{R}\rangle,$ (3.2)
where in between the fourth and last lines we made the substitution $\hat{H}=\hat{\mathbf{P}}^{2}/(2m)+V(\mathbf{R})$. Making use of the momentum-space completeness relationship $\mathds{1}=\int d^{d}p\;|\mathbf{p}\rangle\langle\mathbf{p}|$ and the inner product relationship $\langle\mathbf{R}|\mathbf{p}\rangle=(2\pi\hbar)^{-d/2}\exp[(i/\hbar)\mathbf{p}\cdot\mathbf{R}]$ allows us to recast equation 3.2 as
$C(\mathbf{R},\mathbf{s};\beta)=\frac{g}{(2\pi\hbar)^{d/2}}\int d^{d}p\;\textrm{e}^{-\frac{i}{\hbar}\mathbf{p}\cdot\mathbf{s}}\textrm{e}^{\frac{i}{\hbar}\mathbf{p}\cdot\mathbf{R}}\langle\mathbf{p}|\textrm{e}^{-\beta\frac{\hat{\mathbf{P}}^{2}}{2m}-\beta V(\mathbf{R}+\mathbf{s}/2)}|\mathbf{R}\rangle.$ (3.3)
At this point, we can perform our symmetric semiclassical WKODM expansion following the same techniques used for the usual non-symmetric WKODM method within reference [10]. The process starts by making the substitution $\langle\mathbf{p}|\textrm{Exp}[-\beta\frac{\hat{\mathbf{P}}^{2}}{2m}-\beta V(\mathbf{R}+\mathbf{s}/2)]|\mathbf{R}\rangle=\exp[-\beta H_{cl}(\mathbf{R}+\mathbf{s}/2,\mathbf{p})]\langle\mathbf{p}|\mathbf{R}\rangle\tilde{w}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)$, yielding
$C(\mathbf{R},\mathbf{s};\beta)=\frac{g}{(2\pi\hbar)^{d}}\int d^{d}p\;\textrm{e}^{-\frac{i}{\hbar}\mathbf{p}\cdot\mathbf{s}}\textrm{e}^{-\beta H_{cl}(\mathbf{R}+\mathbf{s}/2,\mathbf{p})}\tilde{w}(\mathbf{R},\mathbf{s},\mathbf{p};\beta),$ (3.4)
where $\tilde{w}$ is a function which obeys
$\textrm{e}^{-\beta H_{cl}(\mathbf{R}+\mathbf{s}/2)-(i/\hbar)\mathbf{p}\cdot\mathbf{R}}\tilde{w}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)=\tilde{u}(\mathbf{R},\mathbf{s},\mathbf{p};\beta),$ (3.5)
and $\tilde{u}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)$ is a function that obeys the Bloch equation [10]:
$\frac{\partial\tilde{u}}{\partial\beta}+\hat{H}\tilde{u}=0.$ (3.6)
For simplicity, we sometimes suppress the $(\mathbf{R},\mathbf{s},\mathbf{p};\beta)$ dependence of both the $\tilde{w}$ and $\tilde{u}$ functions, in a similar form to equation 3.6.
3.2 Symmetric WKODM expansion
Substituting equation 3.5 into equation 3.6, making the second order semiclassical approximation $\tilde{w}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)\approx 1+\hbar\tilde{w}_{1}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)+\hbar^{2}\tilde{w}_{2}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)$, and collecting terms based on their order of $\hbar$ yields
$\hbar\bigg{(}-\frac{i\beta}{m}\mathbf{p}\cdot\nabla_{\mathbf{R}}V(\mathbf{R}+\mathbf{s}/2)+\frac{\partial\tilde{w}_{1}}{\partial\beta}\bigg{)}+\hbar^{2}\bigg{(}-\frac{i\beta}{m}\mathbf{p}\cdot\nabla_{\mathbf{R}}V(\mathbf{R}+\mathbf{s}/2)\tilde{w}_{1}$
$-\frac{\beta^{2}}{2m}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R}+\mathbf{s}/2)\big{)}^{2}+\frac{\beta}{2m}\nabla_{\mathbf{R}}^{2}V(\mathbf{R}+\mathbf{s}/2)+\frac{i}{\hbar}(\mathbf{p}\cdot\nabla_{\mathbf{R}})\tilde{w}_{1}+\tilde{w}_{2}\bigg{)}=0.$ (3.7)
In order for this equation to be true, the expressions within the large brackets multiplying both $\hbar$ and 2superscriptPlanck-constant-over-$\hbar^{2}$ must both be equal to zero. The expression multiplying $\hbar$ can be easily be solved to obtain $\tilde{w}_{1}$, which can then be substituted into the expression multiplying 2superscriptPlanck-constant-over-$\hbar^{2}$ to solve for $\tilde{w}_{2}$. Doing this yields
$\tilde{w}_{1}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)$ $=\bigg{(}\frac{i\beta^{2}}{2m}\bigg{)}\big{(}\mathbf{p}\cdot\nabla_{\mathbf{R}}V(\mathbf{R}+\mathbf{s}/2)\big{)},$ (3.8)
$\tilde{w}_{2}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)$ $=-\frac{\beta^{3}}{4m}\nabla_{\mathbf{R}}^{2}V(\mathbf{R}+\mathbf{s}/2)+\frac{\beta^{3}}{6m}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R}+\mathbf{s}/2)\big{)}^{2}$
$\quad+\frac{\beta^{3}}{6m^{2}}(\mathbf{p}\cdot\nabla_{\mathbf{R}})^{2}V(\mathbf{R}+\mathbf{s}/2)-\frac{\beta^{4}}{8m^{2}}\big{(}\mathbf{p}\cdot\nabla_{\mathbf{R}}V(\mathbf{R}+\mathbf{s}/2)\big{)}^{2}.$ (3.9)
Aside from a minus sign difference in equation 3.8, equations 3.8 and 3.9 are the same in form as their non-symmetric counterparts $w_{1}(\mathbf{r},\mathbf{r}^{\prime};\beta)$ and $w_{2}(\mathbf{r},\mathbf{r}^{\prime};\beta)$ respectively, as shown in A by equations 1.4 and 1.5 respectively [10]. To keep our notation consistent with the non-symmetric WKODM method, we utilize the fact that equation 3.4 remains invariant when the integrand undergoes the momentum transformation $\mathbf{p}\rightarrow-\mathbf{p}$ (this transformation does not include changing the sign of the differential element $d^{d}p$). This is shown below by equation 3.10, where $f(\mathbf{p})$ is any function which is either a polynomial or can expressed as a polynomial by a Taylor expansion.
$\int d^{d}p\;\textrm{e}^{-\frac{i}{\hbar}\mathbf{p}\cdot\mathbf{s}-\beta\frac{\mathbf{p}^{2}}{2m}}f(\mathbf{p})=\int d^{d}p\;\textrm{e}^{-\frac{i}{\hbar}(-\mathbf{p})\cdot\mathbf{s}-\beta\frac{(-\mathbf{p})^{2}}{2m}}f(-\mathbf{p})$ (3.10)
Applying equation 3.10 allows us to re-express equation 3.4 as
$C(\mathbf{R},\mathbf{s};\beta)$ $=\frac{g}{(2\pi\hbar)^{d}}\int d^{d}p\;\textrm{e}^{\frac{i}{\hbar}\mathbf{p}\cdot\mathbf{s}}\textrm{e}^{-\beta H_{cl}(\mathbf{R}+\mathbf{s}/2,\mathbf{p})}\tilde{w}(\mathbf{R},\mathbf{s},-\mathbf{p};\beta)$
$=\frac{g}{(2\pi\hbar)^{d}}\int d^{d}p\;\textrm{e}^{\frac{i}{\hbar}\mathbf{p}\cdot\mathbf{s}}\textrm{e}^{-\beta H_{cl}(\mathbf{R}+\mathbf{s}/2,\mathbf{p})}w(\mathbf{R},\mathbf{s},\mathbf{p};\beta),$ (3.11)
where we have defined $w(\mathbf{R},\mathbf{s},\mathbf{p};\beta)=\tilde{w}(\mathbf{R},\mathbf{s},-\mathbf{p};\beta)$ so that $w(\mathbf{R},\mathbf{s},\mathbf{p};\beta)\approx 1+\hbar w_{1}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)+\hbar^{2}w_{2}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)$ with
$w_{1}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)$ $=\tilde{w}_{1}(\mathbf{R},\mathbf{s},-\mathbf{p};\beta)=-\bigg{(}\frac{i\beta^{2}}{2m}\bigg{)}\big{(}\mathbf{p}\cdot\nabla_{\mathbf{R}}V(\mathbf{R}+\mathbf{s}/2)\big{)},$ (3.12)
$w_{2}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)$ $=\tilde{w}_{2}(\mathbf{R},\mathbf{s},-\mathbf{p};\beta)$
$=-\frac{\beta^{3}}{4m}\nabla_{\mathbf{R}}^{2}V(\mathbf{R}+\mathbf{s}/2)+\frac{\beta^{3}}{6m}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R}+\mathbf{s}/2)\big{)}^{2}$
$\quad+\frac{\beta^{3}}{6m^{2}}(\mathbf{p}\cdot\nabla_{\mathbf{R}})^{2}V(\mathbf{R}+\mathbf{s}/2)-\frac{\beta^{4}}{8m^{2}}\big{(}\mathbf{p}\cdot\nabla_{\mathbf{R}}V(\mathbf{R}+\mathbf{s}/2)\big{)}^{2}.$ (3.13)
Equations 3.12 and 3.13 are in the same form as the non-symmetric expressions of $w_{1}(\mathbf{r},\mathbf{r}^{\prime};\beta)$ and $w_{2}(\mathbf{r},\mathbf{r}^{\prime};\beta)$, given by equations 1.4 and 1.5 within A, respectively. Similar to the procedure we followed both there and within Section 2 for the KODM derivation (based on reference [10]), we are finally in a position to derive the symmetric WKODM $\rho_{WK}(\mathbf{R},\mathbf{s})$ by utilizing inverse Laplace transforms and separating terms based on their order of $\hbar$. We find
$\rho_{WK}(\mathbf{R},\mathbf{s})$ $\approx B_{0}+B_{1}+B_{2};$ (3.14)
$B_{0}$ $=\mathcal{L}^{-1}_{\mu}\bigg{\{}\frac{C_{0}(\mathbf{R},\mathbf{s};\beta)}{\beta}\bigg{\}},$ (3.15)
$B_{1}$ $=\mathcal{L}^{-1}_{\mu}\bigg{\{}\frac{C_{1}(\mathbf{R},\mathbf{s};\beta)}{\beta}\bigg{\}},$ (3.16)
$B_{2}$ $=\mathcal{L}^{-1}_{\mu}\bigg{\{}\frac{C_{2}(\mathbf{R},\mathbf{s};\beta)}{\beta}\bigg{\}},$ (3.17)
where each term of the approximate symmetric Bloch matrix $C(\mathbf{R},\mathbf{s};\beta)\approx C_{0}(\mathbf{R},\mathbf{s};\beta)+C_{1}(\mathbf{R},\mathbf{s};\beta)+C_{2}(\mathbf{R},\mathbf{s};\beta)$ is given by equations 3.18, 3.19 and 3.20. These expressions are in the same form as those of the non-symmetric case shown in A by equations 1.1, 1.2 and 1.3.
$C_{0}(\mathbf{R},\mathbf{s};\beta)$ $=\frac{g}{(2\pi\hbar)^{d}}\int d^{d}p\;\textrm{e}^{-\beta H_{cl}(\mathbf{R}+\mathbf{s}/2,\mathbf{p})+\frac{i}{\hbar}\mathbf{p}\cdot\mathbf{s}},$ (3.18)
$C_{1}(\mathbf{R},\mathbf{s};\beta)$ $=-\frac{ig\hbar\beta^{2}}{2m(2\pi\hbar)^{d}}\int d^{d}p\;\textrm{e}^{-\beta H_{cl}(\mathbf{R}+\mathbf{s}/2,\mathbf{p})+\frac{i}{\hbar}\mathbf{p}\cdot\mathbf{s}}w_{1}(\mathbf{R},\mathbf{s},\mathbf{p};\beta),$ (3.19)
$C_{2}(\mathbf{R},\mathbf{s};\beta)$ $=\frac{g\hbar^{2}}{(2\pi\hbar)^{d}}\int d^{d}p\;\textrm{e}^{-\beta H_{cl}(\mathbf{R}+\mathbf{s}/2,\mathbf{p})+\frac{i}{\hbar}\mathbf{p}\cdot\mathbf{s}}w_{2}(\mathbf{R},\mathbf{s},\mathbf{p};\beta).$ (3.20)
From hereon, the calculations and approximations we perform to fully simplify and combine equations 3.12, 3.13, 3.14, 3.15, 3.16, 3.17, 3.18, 3.19 and 3.20 are identical to those used in Section 2 for the KODM method, hence to avoid redundancy, we will not explicitly show each step. After simplification one exactly finds that in symmetric coordinates $B_{0}=A_{0}$ [equation 2.17], $B_{1}=A_{1}$ [equation 2.23], and $B_{2}=A_{2}$ [equation 2.24]. Therefore, combining these results together via equation 3.14, we find that the symmetric WKODM is exactly identical to the symmetric KODM [equation 2.25], which we know is exactly the $d\geq 1$-dimensional GVODM originally proposed in reference [1].
4 Concluding remarks
In summary we have studied the approximate semiclassical ODM of a zero temperature non-interacting Fermi gas obtained through three different gradient expansion techniques: the KODM, WKODM, and GVODM methods. We have resolved the inconsistencies discussed in the literature [1, 21, 22], and have shown that, regardless of whichever method is used, the ODMs are Hermitian, idempotent, and identical when cast into symmetric coordinates $(\mathbf{R},\mathbf{s})$. Inconsistencies only arise when the semiclassical expansions are truncated in terms of non-symmetric coordinates $(\mathbf{r},\mathbf{r}^{\prime})$, which occur with the non-symmetric KODM [equation 2.1] and WKODM methods [A]. Additionally, we have provided two separate and explicit derivations of the $d\geq 1$-dimensional GVODM originally proposed by reference [1] based on their $d=1,2,3,4$ expressions. The first derivation (Section 2) was accomplished by transforming the known $d\geq 1$-dimensional expression of the KODM [20] into symmetric coordinates, while the second derivation (Section 3) transformed the known $d\geq 1$-dimensional WKODM into symmetric coordinates from the very beginning. For simplicity, we only dealt with semiclassical terms up to second order in $\hbar$, nevertheless, our results still hold when higher order terms in the ODMs are accounted for.
L.M.F., D.E., and P.C. dedicate this paper to the memory of their late supervisor, Brandon van Zyl, who passed away in 2020. This work originally began as an undergraduate honours thesis research project between the authors during the summer of 2017. The goal was to follow up on comments in reference [22], and address the subtle and consistently overlooked gap in the literature (see references/comments 11 and 12 in reference [22]). We decided to continue where we left off, and finally completed this work to publish it in honour of B.V.Z., which was the original intention before his untimely passing. This work would not have been possible without the contribution of B.V.Z. The authors would like to thank the Natural Sciences and Engineering Research Council of Canada (NSERC) for funding during the initial stages of this work in 2017. We would also like to thank Karl-Peter Marzlin, Duncan O’Dell, and Wyatt Kirkby for useful discussions.
Appendix A Non-symmetric WKODM
In this section, we apply the WKODM method in non-symmetric coordinates, and obtain an expression identical to the non-symmetric KODM of equation 2.1, finding that $\rho_{WK}(\mathbf{r},\mathbf{r}^{\prime})=\rho_{K}(\mathbf{r},\mathbf{r}^{\prime})$. This is not to be confused with our symmetric WKODM method derivation in Section 3, where we similarly show that $\rho_{WK}(\mathbf{R},\mathbf{s})=\rho_{GV}(\mathbf{R},\mathbf{s})$.
In showing $\rho_{WK}(\mathbf{r},\mathbf{r}^{\prime})=\rho_{K}(\mathbf{r},\mathbf{r}^{\prime})$, we see that equations 2.2, 2.3 and 2.4 can alternatively be expressed in terms of inverse Laplace transformations via equations 2.5, 2.6 and 2.7. Using this fact we are able to show in Section 2 how to properly transform the KODM into symmetric coordinates $(\mathbf{R},\mathbf{s})$ by performing the transformation of variables before the inverse Laplace transforms are taken. Making the transformation after would instead result in additional terms being incorrectly unaccounted for.
In the non-symmetric WKODM method the $d\geq 1$-dimensional Bloch matrix is approximated to second order in $\hbar$ by $C(\mathbf{r},\mathbf{r}^{\prime};\beta)\approx C_{0}(\mathbf{r},\mathbf{r}^{\prime};\beta)+C_{1}(\mathbf{r},\mathbf{r}^{\prime};\beta)+C_{2}(\mathbf{r},\mathbf{r}^{\prime};\beta)$ where
$C_{0}(\mathbf{r},\mathbf{r}^{\prime};\beta)$ $=\frac{g}{(2\pi\hbar)^{d}}\int d^{d}p\;\textrm{e}^{-\beta H_{cl}(\mathbf{r},\mathbf{p})+\frac{i}{\hbar}\mathbf{p}\cdot\mathbf{s}},$ (1.1)
$C_{1}(\mathbf{r},\mathbf{r}^{\prime};\beta)$ $=-\frac{ig\hbar\beta^{2}}{2m(2\pi\hbar)^{d}}\int d^{d}p\;\textrm{e}^{-\beta H_{cl}(\mathbf{r},\mathbf{p})+\frac{i}{\hbar}\mathbf{p}\cdot\mathbf{s}}w_{1}(\mathbf{r},\mathbf{p};\beta),$ (1.2)
$C_{2}(\mathbf{r},\mathbf{r}^{\prime};\beta)$ $=\frac{g\hbar^{2}}{(2\pi\hbar)^{d}}\int d^{d}p\;\textrm{e}^{-\beta H_{cl}(\mathbf{r},\mathbf{p})+\frac{i}{\hbar}\mathbf{p}\cdot\mathbf{s}}w_{2}(\mathbf{r},\mathbf{p};\beta),$ (1.3)
and
$w_{1}(\mathbf{r},\mathbf{p};\beta)$ $=-\bigg{(}\frac{i\beta^{2}}{2m}\bigg{)}\big{(}\mathbf{p}\cdot\nabla_{\mathbf{r}}V(\mathbf{r})\big{)},$ (1.4)
$w_{2}(\mathbf{r},\mathbf{p};\beta)$ $=-\frac{\beta^{3}}{4m}\nabla_{\mathbf{r}}^{2}V(\mathbf{r})+\frac{\beta^{3}}{6m}\big{(}\nabla_{\mathbf{r}}V(\mathbf{r})\big{)}^{2}$
$\quad+\frac{\beta^{3}}{6m^{2}}(\mathbf{p}\cdot\nabla_{\mathbf{r}})^{2}V(\mathbf{r})-\frac{\beta^{4}}{8m^{2}}\big{(}\mathbf{p}\cdot\nabla_{\mathbf{r}}V(\mathbf{r})\big{)}^{2}.$ (1.5)
$H_{cl}(\mathbf{r},\mathbf{p})=\mathbf{p}^{2}/(2m)+V(\mathbf{r})$ and the details of where these expressions come from are described in reference [10], although they focus on the specific case of a diagonal system where $\mathbf{r}=\mathbf{r}^{\prime}$ so that $\mathbf{s}=0$. Our case is more general since we assume a small but finite $\mathbf{s}$. Our symmetric WKODM derivation in Section 3 will also result in similar expressions, but in terms of symmetric coordinates.
Combining equations 1.1, 1.2, 1.3, 1.4 and 1.5 is precisely the origin of equations 2.8, 2.9 and 2.10. In order to justify expressing the KODM in terms of inverse Laplace transforms we show below that in non-symmetric coordinates the KODM [equations 2.1, 2.2, 2.3 and 2.4] is exactly obtained by substituting equations 2.8, 2.9 and 2.10 into equations 2.5, 2.6 and 2.7, respectively. Starting with the $A_{0}$ term, equations 2.5 and 2.8 yield
$A_{0}=\mathcal{L}^{-1}_{\mu}\bigg{\{}\frac{C_{0}(\mathbf{r},\mathbf{r}^{\prime};\beta)}{\beta}\bigg{\}}$ $=g\bigg{(}\frac{m}{2\pi\hbar^{2}}\bigg{)}^{d/2}\mathcal{L}_{\mu}^{-1}\bigg{\{}\beta^{-d/2-1}\textrm{e}^{-\frac{m\mathbf{s}^{2}}{2\hbar^{2}\beta}-\beta V(\mathbf{r})}\bigg{\}}$
$=\frac{gk_{F}^{d}}{(2\pi z)^{d/2}}J_{\frac{d}{2}}(z).$
Similarly for $A_{1}$, equations 2.6 and 2.9 give
$A_{1}=\mathcal{L}^{-1}_{\mu}\bigg{\{}\frac{C_{1}(\mathbf{r},\mathbf{r}^{\prime};\beta)}{\beta}\bigg{\}}$ $=-\frac{gm^{d/2-1}\hbar^{2-d}}{2(2\pi)^{d/2}}\big{(}\nabla_{\mathbf{r}}V(\mathbf{r})\cdot\nabla_{\mathbf{s}}\big{)}\mathcal{L}_{\mu}^{-1}\bigg{\{}\beta^{-d/2+1}\textrm{e}^{-\frac{m\mathbf{s}^{2}}{2\hbar^{2}\beta}-\beta V(\mathbf{r})}\bigg{\}}$
$=-\frac{gk_{F}^{d-3}z^{2}}{4(2\pi z)^{d/2}}J_{\frac{d}{2}-1}(z)\bigg{(}\nabla_{\mathbf{r}}k_{F}^{2}\cdot\frac{\mathbf{s}}{s}\bigg{)}.$
For $A_{2}$, combing equations 2.7 and 2.10 returns
$A_{2}=\mathcal{L}^{-1}_{\mu}\bigg{\{}\frac{C_{2}(\mathbf{r},\mathbf{r}^{\prime};\beta)}{\beta}\bigg{\}}$ $=-\frac{gm^{d/2-1}}{12(2\pi)^{d/2}\hbar^{d-2}}\big{(}\nabla_{\mathbf{r}}^{2}V(\mathbf{r})\big{)}\mathcal{L}_{\mu}^{-1}\bigg{\{}\beta^{-d/2+1}\textrm{e}^{-\frac{m\mathbf{s}^{2}}{2\hbar^{2}\beta}-\beta V(\mathbf{r})}\bigg{\}}$
$\quad+\frac{gm^{d/2-1}}{24(2\pi)^{d/2}\hbar^{d-2}}\big{(}\nabla_{\mathbf{r}}V(\mathbf{r})\big{)}^{2}\mathcal{L}_{\mu}^{-1}\bigg{\{}\beta^{-d/2+2}\textrm{e}^{-\frac{m\mathbf{s}^{2}}{2\hbar^{2}\beta}-\beta V(\mathbf{r})}\bigg{\}}$
$\quad+\frac{gm^{d/2}}{8(2\pi)^{d/2}\hbar^{d}}\big{(}\nabla_{\mathbf{r}}V(\mathbf{r})\cdot{\mathbf{s}}\big{)}\mathcal{L}_{\mu}^{-1}\bigg{\{}\beta^{-d/2+1}\textrm{e}^{-\frac{m\mathbf{s}^{2}}{2\hbar^{2}\beta}-\beta V(\mathbf{r})}\bigg{\}}$
$\quad-\frac{gm^{d/2}}{6(2\pi)^{d/2}\hbar^{d}}\Big{(}\nabla_{\mathbf{r}}\big{(}\nabla_{\mathbf{r}}V(\mathbf{r})\cdot{\mathbf{s}}\big{)}\cdot{\mathbf{s}}\Big{)}\mathcal{L}_{\mu}^{-1}\bigg{\{}\beta^{-d/2}\textrm{e}^{-\frac{m\mathbf{s}^{2}}{2\hbar^{2}\beta}-\beta V(\mathbf{r})}\bigg{\}}$
$\implies A_{2}$ $=\frac{gk_{F}^{d-4}z^{2}}{24(2\pi z)^{d/2}}J_{\frac{d}{2}-2}(z)\big{(}\nabla_{\mathbf{r}}^{2}k_{F}^{2}\big{)}+\frac{gk_{F}^{d-6}z^{3}}{96(2\pi z)^{d/2}}J_{\frac{d}{2}-3}(z)\big{(}\nabla_{\mathbf{r}}k_{F}^{2}\big{)}^{2}$
$\quad+\frac{gk_{F}^{d-6}z^{4}}{32(2\pi z)^{d/2}}J_{\frac{d}{2}-2}(z)\bigg{(}\nabla_{\mathbf{r}}k_{F}^{2}\cdot\frac{\mathbf{s}}{s}\bigg{)}^{2}+\frac{gk_{F}^{d-4}z^{3}}{12(2\pi z)^{d/2}}J_{\frac{d}{2}-1}(z)\bigg{[}\nabla_{\mathbf{r}}\bigg{(}\nabla_{\mathbf{r}}k_{F}^{2}\cdot\frac{\mathbf{s}}{s}\bigg{)}\cdot\frac{\mathbf{s}}{s}\bigg{]}.$
Each term was computed in terms of non-symmetric coordinates so that $k_{F}=k_{F}(\mathbf{r})$, and the inverse Laplace transform identity equation 2.4 was used in each case (with $\mathbf{R}$ replaced by $\mathbf{r}$). Additionally for the $A_{2}$ term, two $d\geq 1$-dimensional Gaussian integrals similar to that in equation 2.8 were solved straightforwardly, and lastly the vector identity $(\mathbf{s}\cdot\nabla_{\mathbf{r}})^{2}V(\mathbf{r})=\nabla_{\mathbf{r}}\big{(}\nabla_{\mathbf{r}}V(\mathbf{r})\cdot\mathbf{s}\big{)}\cdot\mathbf{s}$ was applied.
Appendix B Useful relationships
B.1 Bessel function identities
In this subsection, we list specific identities of cylindrical Bessel functions $J_{\nu}(z)$ of order $\nu$ which are used throughout our work [25].
$J_{\nu}(z)$ $=\frac{2(\nu+1)}{z}J_{\nu+1}(z)-J_{\nu+2}(z)$ (2.1)
$J_{\nu}(z)$ $=\frac{\big{(}4(\nu+1)(\nu+2)-z^{2}\big{)}J_{\nu+2}(z)-2z(\nu+1)J_{\nu+3}(z)}{z^{2}}$ (2.2)
$\frac{\partial(z^{-\nu}J_{\nu}(z))}{\partial z}$ $=-z^{-\nu}J_{\nu+1}(z)$ (2.3)
B.2 Inverse Laplace transform identities
In this subsection, we evaluate the following inverse Laplace transformation which will be used in various places throughout our work.
$\mathcal{L}_{\mu}^{-1}\bigg{\{}\beta^{-d/2-\xi}\textrm{e}^{-\frac{m\mathbf{s}^{2}}{2\hbar^{2}\beta}-\beta V(\mathbf{R})}\bigg{\}}$ $=\int_{0}^{\mu}\bigg{(}\frac{2\hbar^{2}(\mu-\mu^{\prime})}{m\mathbf{s}^{2}}\bigg{)}^{(d/2+\xi-1)/2}$
×$\quad\times J_{\frac{d}{2}+\xi-1}\Bigg{(}\sqrt{\frac{2m\mathbf{s}^{2}}{\hbar^{2}}(\mu-\mu^{\prime})}\Bigg{)}\delta\big{(}\mu^{\prime}-V(\mathbf{R})\big{)}d\mu^{\prime}$
$=\bigg{(}\frac{2\hbar^{2}\big{(}\mu-V(\mathbf{R})\big{)}}{m\mathbf{s}^{2}}\bigg{)}^{(d/2+\xi-1)/2}$
×$\quad\times J_{\frac{d}{2}+\xi-1}\Bigg{(}\sqrt{\frac{2m\mathbf{s}^{2}}{\hbar^{2}}\big{(}\mu-V(\mathbf{R})\big{)}}\Bigg{)}$
$=\Bigg{(}\frac{\hbar^{4}k_{F}^{2}}{m^{2}\mathbf{s}^{2}}\Bigg{)}^{(d/2+\xi-1)/2}J_{\frac{d}{2}+\xi-1}(z),$ (2.4)
where $\xi$ is some real constant which, in this paper, only ever takes values $\xi=1,0,-1,-2$, while $\beta$ [not to be confused with the inverse of the temperature $(k_{B}T)^{-1}$] and the chemical potential $\mu$ are conjugate variables of the Laplace transformation. Going between the first and second lines in the derivation of equation 2.4, we utilized three common inverse Laplace transforms which are listed below [10, 25], and going between the third and fourth lines we made the substitution $\mu-V(\mathbf{R})=\frac{\hbar^{2}}{2m}k_{F}(\mathbf{R})^{2}$, suppressed the $\mathbf{R}$ dependent notation of $k_{F}$ so that $k_{F}=k_{F}(\mathbf{R})$, and defined $z=k_{F}s$.
$\mathcal{L}_{\mu}^{-1}\big{\{}\textrm{e}^{-\beta V(\mathbf{R})}\big{\}}$ $=\delta\big{(}\mu-V(\mathbf{R})\big{)},$
[…]
$\mathcal{L}_{\mu}^{-1}\bigg{\{}\frac{1}{\beta^{d/2+\xi}}\textrm{e}^{-\frac{m\mathbf{s}^{2}}{2\hbar^{2}\beta}}\bigg{\}}$ $=\bigg{(}\frac{2\hbar^{2}\mu}{m\mathbf{s}^{2}}\bigg{)}^{(d/2+\xi-1)/2}J_{\frac{d}{2}+\xi-1}\Bigg{(}\sqrt{\frac{2m\mathbf{s}^{2}}{\hbar^{2}}\mu}\Bigg{)}.$
$F(\beta)$ and $G(\beta)$ are arbitrary functions representing the Laplace transforms of the functions $f(\mu)$ and $g(\mu)$, respectively. The third of the above three common inverse Laplace transform only holds true when $\frac{m\mathbf{s}^{2}}{2\hbar^{2}}>-1$. In our case, this is always true since this combination of values is in fact always greater than 00. We also emphasize that equation 2.4 is utilized in A with $\mathbf{R}$ replaced by $\mathbf{r}$.
B.3 Gradient identities
In this subsection we list some exact and approximate identities between both gradients and derivatives with respect to $\mathbf{r}$, $\mathbf{R}$, $\mathbf{s}$, and $z$. These are applied in various places throughout our work and are given by the following expressions:
$\nabla_{\mathbf{r}}$ $=\frac{1}{2}\nabla_{\mathbf{R}}+\nabla_{\mathbf{s}},$ (2.5)
$\nabla_{\mathbf{r}}^{2}$ $=\frac{1}{4}\nabla_{\mathbf{R}}^{2}+\nabla_{\mathbf{s}}^{2}+\nabla_{\mathbf{R}}\cdot\nabla_{\mathbf{s}},$ (2.6)
$\nabla_{\mathbf{s}}$ $=k_{F}\frac{\mathbf{s}}{s}\frac{\partial}{\partial z},$ (2.7)
$\nabla_{\mathbf{r}}V(\mathbf{r})$ $\approx\nabla_{\mathbf{R}}V(\mathbf{R})+\frac{1}{2}\nabla_{\mathbf{R}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)},$ (2.8)
$\nabla_{\mathbf{r}}^{2}V(\mathbf{r})$ $\approx\nabla_{\mathbf{R}}^{2}V(\mathbf{R}),$ (2.9)
$\nabla_{\mathbf{r}}\bigg{(}\nabla_{\mathbf{r}}V(\mathbf{r})\cdot\frac{\mathbf{s}}{s}\bigg{)}\cdot\frac{\mathbf{s}}{s}$ $=\nabla_{\mathbf{R}}\bigg{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\frac{\mathbf{s}}{s}\bigg{)}\cdot\frac{\mathbf{s}}{s}.$ (2.10)
As is the case throughout the rest of the paper, an equal ($=$) sign denotes an exact identity, and an approximation sign ($\approx$) denotes an approximate one, where only terms up to and including second order in $\mathbf{s}$ (equivalently $\nabla_{\mathbf{R}}$ and $\hbar$) are retained. Equations 2.5, 2.6 and 2.7 are straightforward to show via equation 1.3 and the definition of $z$, but equations 2.8, 2.9 and 2.10 are more complicated and deserve their own proofs. Starting with equation 2.8 we show that
$\nabla_{\mathbf{r}}V(\mathbf{r})=\nabla_{\mathbf{r}}V(\mathbf{R}+\mathbf{s}/2)$ $\approx\nabla_{\mathbf{r}}V(\mathbf{R})+\frac{1}{2}\nabla_{\mathbf{r}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}+\frac{1}{8}\nabla_{\mathbf{r}}\Big{[}\nabla_{\mathbf{R}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}\cdot\mathbf{s}\Big{]}$
$\approx\nabla_{\mathbf{R}}V(\mathbf{R})+\frac{1}{4}\nabla_{\mathbf{R}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}+\frac{1}{8}\nabla_{\mathbf{s}}\Big{[}\nabla_{\mathbf{R}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}\cdot\mathbf{s}\Big{]}$
$=\nabla_{\mathbf{R}}V(\mathbf{R})+\frac{1}{4}\nabla_{\mathbf{R}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}+\frac{1}{4}\nabla_{\mathbf{R}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}$
$=\nabla_{\mathbf{R}}V(\mathbf{R})+\frac{1}{2}\nabla_{\mathbf{R}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)},$
where equation 2.12 was applied in the first line and equation 2.5 between lines one and two. For equation 2.9, we find
$\nabla_{\mathbf{r}}^{2}V(\mathbf{r})$ $=\nabla_{\mathbf{r}}^{2}V(\mathbf{R}+\mathbf{s}/2)$
$\approx\nabla_{\mathbf{r}}^{2}V(\mathbf{R})+\frac{1}{2}\nabla_{\mathbf{r}}^{2}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}+\frac{1}{8}\nabla_{\mathbf{r}}^{2}\Big{[}\nabla_{\mathbf{R}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}\cdot\mathbf{s}\Big{]}$
$\approx\frac{1}{4}\nabla_{\mathbf{R}}^{2}V(\mathbf{R})+\frac{1}{8}\nabla_{\mathbf{s}}^{2}\Big{[}\nabla_{\mathbf{R}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}\cdot\mathbf{s}\Big{]}+\frac{1}{2}\big{(}\nabla_{\mathbf{R}}\cdot\nabla_{\mathbf{s}}\big{)}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}$
$=\frac{3}{4}\nabla_{\mathbf{R}}^{2}V(\mathbf{R})+\frac{1}{4}\nabla_{\mathbf{s}}^{2}\Big{[}\nabla_{\mathbf{R}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}\cdot\mathbf{s}\Big{]}$
$=\frac{3}{4}\nabla_{\mathbf{R}}^{2}V(\mathbf{R})+\frac{1}{4}\nabla_{\mathbf{R}}^{2}V(\mathbf{R})$
$=\nabla_{\mathbf{R}}^{2}V(\mathbf{R}),$
where equation 2.12 was applied in the first line and equation 2.6 in between lines one and two. For equation 2.10, it is useful write gradients and their dot products in terms of sums and products of vector components. We find the following
$\nabla_{\mathbf{r}}\bigg{(}\nabla_{\mathbf{r}}V(\mathbf{r})\cdot\frac{\mathbf{s}}{s}\bigg{)}\cdot\frac{\mathbf{s}}{s}$ $=\frac{1}{s}\sum_{i=1}^{d}\sum_{j=1}^{d}s_{i}\frac{\partial}{\partial r_{i}}\bigg{[}\frac{1}{s}s_{j}\bigg{(}\frac{1}{2}\frac{\partial V(\mathbf{R}+\mathbf{s}/2)}{\partial R_{j}}+\frac{\partial V(\mathbf{R}+\mathbf{s}/2)}{\partial s_{j}}\bigg{)}\Bigg{]}$
$=\frac{1}{s^{2}}\sum_{i=1}^{d}\sum_{j=1}^{d}s_{i}\Bigg{[}\frac{1}{2}\bigg{(}\delta_{ij}\frac{\partial V(\mathbf{R}+\mathbf{s}/2)}{\partial R_{j}}+\frac{s_{j}}{2}\frac{\partial^{2}V(\mathbf{R}+\mathbf{s}/2)}{\partial R_{i}\partial R_{j}}$
$\quad+s_{j}\frac{\partial^{2}V(\mathbf{R}+\mathbf{s}/2)}{\partial s_{i}\partial R_{j}}-\frac{s_{i}s_{j}}{s^{2}}\frac{\partial V(\mathbf{R}+\mathbf{s}/2)}{\partial R_{j}}\bigg{)}$
$\quad+\bigg{(}\delta_{ij}\frac{\partial V(\mathbf{R}+\mathbf{s}/2)}{\partial s_{j}}+\frac{s_{j}}{2}\frac{\partial^{2}V(\mathbf{R}+\mathbf{s}/2)}{\partial R_{i}\partial s_{j}}$
$\quad+s_{j}\frac{\partial^{2}V(\mathbf{R}+\mathbf{s}/2)}{\partial s_{i}\partial s_{j}}-\frac{s_{i}s_{j}}{s^{2}}\frac{\partial V(\mathbf{R}+\mathbf{s}/2)}{\partial s_{j}}\bigg{)}\Bigg{]}$
$=\frac{1}{s^{2}}\sum_{i=1}^{d}\sum_{j=1}^{d}\bigg{(}\delta_{ij}\frac{\partial V(\mathbf{R})}{\partial R_{j}}+s_{j}\frac{\partial^{2}V(\mathbf{R})}{\partial R_{i}\partial R_{j}}-\frac{s_{i}s_{j}}{s^{2}}\frac{\partial V(\mathbf{R})}{\partial R_{j}}\bigg{)}$
$=\frac{1}{s^{2}}\sum_{i=1}^{d}\sum_{j=1}^{d}s_{i}s_{j}\frac{\partial^{2}V(\mathbf{R})}{\partial R_{i}\partial R_{j}}$
$=\nabla_{\mathbf{R}}\bigg{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\frac{\mathbf{s}}{s}\bigg{)}\cdot\frac{\mathbf{s}}{s},$
where we applied equation 2.5 in line one, and $\delta_{ij}$ is a Kronecker delta: $\delta_{ij}=1$ when $i=j$ and $\delta_{ij}=0$ when $i\neq j$. In between lines two and four we used the fact that both $\partial V(\mathbf{R}+\mathbf{s}/2)/\partial R_{i}=\big{(}\partial(R_{i}+s_{i}/2)/\partial R_{i}\big{)}\partial V(\mathbf{R}+\mathbf{s}/2)/\partial(R_{i}+s_{i}/2)=\partial V(\mathbf{R})/\partial R_{i}$ and $\partial V(\mathbf{R}+\mathbf{s}/2)/\partial s_{i}=\big{(}\partial(R_{i}+s_{i}/2)/\partial s_{i}\big{)}\partial V(\mathbf{R}+\mathbf{s}/2)/\partial(R_{i}+s_{i}/2)=(1/2)\partial V(\mathbf{R})/\partial R_{i}$, where $i=1,2,...,d$ and $s=\sqrt{\mathbf{s}\cdot\mathbf{s}}$ with $\mathbf{s}=\mathbf{r}-\mathbf{r^{\prime}}$.